OpenStudy (anonymous):

Whats next? Please help. y - b = mx + b - b we subtract b from each side. y - b = mx We simplify Now what do we do next? I am lost.. Do we... A. Add x to each side B. Multiply each side by x C. Divide each side by m D. Subtract b from each side Please answer! Thanks. :)

OpenStudy (texaschic101):

are you trying to get it to y = mx + b ??

OpenStudy (texaschic101):

what are you trying to do ?

OpenStudy (anonymous):

Yes! @texaschic101

OpenStudy (anonymous):

y = mx + b @texaschic101

OpenStudy (anonymous):

I believe we divide each side by m, but I am unsure.

OpenStudy (texaschic101):

y - b = mx to get y by itself on one side, you would have to add b to both sides But that is not one of your answer choices. Are you sure your trying to get it to y = mx + b ??

OpenStudy (texaschic101):

what is the original equation you started with ?

OpenStudy (anonymous):

After we simplify y - b = mx what do we do? Thats where I am stuck. The equation is y - b = mx + b - b. @texaschic101

OpenStudy (texaschic101):

y - b = mx add b to both sides y - b + b = mx + b combine like terms y = mx + b That should be the answer....I don't understand why it isn't an answer choice.

OpenStudy (anonymous):

I need to identify the next step in solving this equation.. Do you understand? @texaschic101

OpenStudy (texaschic101):

the next step is to add b to both sides

OpenStudy (anonymous):

It only gives me the option D. Subtract b from each side.. :/

OpenStudy (texaschic101):

is the problem written correctly ?

OpenStudy (anonymous):

See, after y - b = mx What is the next step? Either we A. Add x to each side B. Multiply each side by x C. Divide each side by m Or lastly, D. Subtract b from each side.

OpenStudy (texaschic101):

@souvik .....please help

OpenStudy (texaschic101):

if you want y = mx + b and you have y - b = mx, the next step is to add b to each side. Thats all I can say.

OpenStudy (texaschic101):

souvik...this is the only way, isn't it ?

OpenStudy (anonymous):

I'm not trying to get an end result I am just trying to figure out what do I do next. @texaschic101

OpenStudy (souvik):

you have to add b to get to your answer...

OpenStudy (texaschic101):

thats the only way

OpenStudy (anonymous):

okay, I just chose A. Thanks guys. Can you help me solve another? :)

OpenStudy (texaschic101):

lets see it :)

OpenStudy (anonymous):

Identify the first step in solving the equation below. P = 21 + 2w

OpenStudy (texaschic101):

you need to tell your teacher that the answer was not in the answer choices on your last question, because if you put A it will be wrong

OpenStudy (texaschic101):

what do you mean the first step ?

OpenStudy (texaschic101):

are there any numbers to sub in ?

OpenStudy (anonymous):

A. Subtract 2w from each side. B. Multiply each side by w. C. Add 2w to each side. D. Divide each side by 2.

OpenStudy (texaschic101):

what result are you trying to get ?

OpenStudy (anonymous):

It says solve for L in this equation P = 2L + 2w

OpenStudy (texaschic101):

first subtract 2w from each side

OpenStudy (texaschic101):

because you want 2L on one side by itself

OpenStudy (anonymous):

Thank you so much @texaschic101 Um, the first question was to C. Divide each side by m.

OpenStudy (anonymous):

lol... I was right the first time.